OpenStudy (anonymous):

How to solve? I tried KCL with variables but the variables kept canceling out.

OpenStudy (anonymous):

Pic:

OpenStudy (ash2326):

I'll help you, let me mark the voltages

OpenStudy (anonymous):

Thank you.

OpenStudy (ash2326):

Notice where I've taken ground, with respect to it, voltages are marked

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

okay how do u solved

OpenStudy (ash2326):

Now write node equation at node, marked with voltage V, can you write?

OpenStudy (anonymous):

-V+ ((12-V)/2) -4 = 0

OpenStudy (anonymous):

Right?

OpenStudy (ash2326):

A small error, it'd be \[\frac{0-V}{1K}+\frac{12-V}{2K}+4 mA=0\] \[-V+\frac{12-V}{2}+4=0\] Now find V from this

OpenStudy (anonymous):

r u sure ash

OpenStudy (ash2326):

yeah, summing all the incoming currents

OpenStudy (ash2326):

*outgoing

OpenStudy (ash2326):

Omg, I'm doing wrong

OpenStudy (ash2326):

Let me rewrite

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

OK

OpenStudy (ash2326):

this is correct, summing all incoming currents \[-V+\frac{12-V}{2}+4=0\]

OpenStudy (anonymous):

Why don't you subtract 4 mA, since it is incoming.

OpenStudy (anonymous):

*?

OpenStudy (ash2326):

I have all incoming currents

OpenStudy (anonymous):

yeah keep going

OpenStudy (anonymous):

But why is the first term -V? Shouldn't it be positive V since it is incoming?

OpenStudy (ash2326):

 A and B are potentials